Car A and car B are traveling in the same direction. Car A is traveling at a constant velocity of 60 mph and car B is traveling at a constant velocity of 75mph. If car A reaches mile marker M 20 minutes before car B, how much time will elapse between the time that car B reaches mile marker M and the time that car B reaces car A?: Car A and car B are traveling in the same direction. Car A is traveling at a constant velocity of 60 mph and car B is traveling at a constant velocity of 75mph. If car A reaches mile marker M 20 minutes before car B, how much time will elapse between the time that car B reaches mile marker M and the time that car B reaces car A?
Answer by ankor@dixie-net.com(3894)

Car A and car B are traveling in the same direction. Car A is traveling at a constant velocity of 60 mph and car B is traveling at a constant velocity of 75mph. If car A reaches mile marker M 20 minutes before car B, how much time will elapse between the time that car B reaches mile marker M and the time that car B reaches car A?
:
During the 20 min that elapses for car B to get to the marker, car A will have traveled:
20/60 * 60 = 20 miles (from the marker)
:
Let t = time for the car B to catch car A
:
Write a distance equation Distance = speed * time
:
Car B distance = Car A distance + 20 miles
75t = 60t + 20
75t - 60t = 20
15t = 20
t = 20/15
t = = 11/3 hrs
:
Another easy way:
Using relative speed. 75 - 60 = 15 mph
Assume that Car A is stationary and car B is traveling at 15 mph
:
How long will it take Car B to travel 20 mi at 15 mph?
time = 20/15 = 11/3 hrs